# Polaritonic control of blackbody infrared radiative dissociation

### v1

@@ Abstract
Vibrational strong light-matter coupling offers a promising approach for controlling chemical reactivity with infrared microcavities. This study explores the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes. Using a Master equation approach, we simulate the effects of infrared field confinement and vibrational strong coupling on BIRD rates for diatomic molecules. We present a framework explaining how infrared microcavities influence BIRD kinetics, highlighting the importance of overtone transitions in the process. Our findings reveal conditions for significant enhancement and mild suppression of radiative dissociation, establishing upper bounds for BIRD rates under weak and strong coupling. These results provide new strategies and limitations for controlling reactive processes with infrared resonators.
@@ I Introduction
Chemical reaction control is critical in applications ranging from industrial catalysis to medicine. Conventional methods for manipulating reaction rates involve varying temperature, pressure, or catalyst composition. An unconventional approach for controlling thermal reaction kinetics has been recently reported Thomas et al. (2016, 2019); Hirai et al. (2020); Ahn et al. (2023) employing infrared (IR) microcavities Kavokin et al. (2017) consisting of two moderate-quality mirrors separated by a distance $L_{C}$ of the order of IR wavelengths resonant with typical molecular vibrations Nagarajan, Thomas, and Ebbesen (2021). When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device Vinogradov (1992); Shalabney et al. (2015); Simpkins et al. (2015). Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction ratesGalego et al. (2019); Campos-Gonzalez-Angulo, Ribeiro, and Yuen-Zhou (2019); Zhdanov (2020); Wang, Flick, and Yelin (2022); Lindoy, Mandal, and Reichman (2022); Schäfer et al. (2022); Sun and Vendrell (2022); Ahn et al. (2023); Montillo Vega, Ying, and Huo (2024); Moiseyev (2024). Still, several fundamental questions remain Simpkins, Dunkelberger, and Owrutsky (2021); Campos-Gonzalez-Angulo et al. (2023), including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted. In this study, we investigate gas-phase unimolecular dissociation activated by the absorption of microcavity blackbody infrared radiation Dunbar (2004) under weak and strong coupling with a suitable material.
@@
Blackbody infrared radiative dissociation (BIRD) was proposed by Perrin in 1913 as a mechanism for gas phase unimolecular dissociation Perrin (1970). However, experimental observations in the 1920s ruled BIRD out in favor of the collisional activation mechanism Lindemann et al. (1922); Daniels (1928); King and Laidler (1984). Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressuresThoelmann, Tonner, and McMahon (1994); Dunbar et al. (1995); Schnier et al. (1996). Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several simple and complex systems, including ion clusters, transition-metal complexes, and biopolymers Dunbar (2004).
@@
BIRD rates are heavily influenced by the thermal radiation spectrum. Given that free space and microcavity electromagnetic density of states and thermal radiation density can be significantly different Kavokin et al. (2017), microcavity BIRD is expected to proceed with different rates compared to free space BIRD. However, this difference’s order of magnitude, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime. Our study provides answers to these questions. Our quantitative analysis focuses on the dissociation of diatomic molecules induced by multiphoton or multipolariton absorption in microcavities, but our qualitative conclusions are generalizable to polyatomic systems.
@@
The primary aim of this study is to characterize thermal radiative dissociation in microcavities. We employ a Master equation treatment to investigate blackbody infrared radiative dissociation for diatomic molecules in (i) free space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material. Our results provide upper bounds for enhancement and suppression of BIRD in weakly and strongly coupled microcavities and identify overtone transitions as key ingredients for the observed modulation of BIRD. In Sec. II, we describe our methodology, Sec. III provides our main results and related discussion, and Sec. IV summarizes our main results and conclusions.
@@ II Methods
II.1 Master Equation for diatomic BIRD
Our analysis of BIRD in microcavities employs the same tools successfully applied to investigate this process in free space Dunbar (2004); Dunbar et al. (1995); Stevens et al. (2002); Dunbar and McMahon (1998). These include a Markovian Master equation describing the time-evolution of vibrational quantum state populations with energy below the dissociation threshold. Rotational degrees of freedom are ignored since at the temperatures of interest to BIRD, the rotational dynamics is essentially classical, only introducing a nonzero linewidth to the vibrational absorption and emission spectrum. We discuss further implications of this condition in the context of our results below.
The rate of change of the population of the $i$th vibrational state ($N_{i}$) is given by
@@ II.1 Master Equation for diatomic BIRD
where $k_{ij}$ denotes rate of absorption when $i>j$ and emission when $i<j$ and $i_{\text{max}}$ refers to the highest energy bound state (Fig. 1a). We assume that the state $i_{\text{max}}$ undergoes irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$. For the sake of simplicity, we have assumed the thermalization of the background radiation is faster than all considered radiative transitions. Equation 1 can be written in matrix notation as
@@
where N is a column vector containing the vibrational populations and $\mathbf{J}$ is the transition matrix. The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction Valance and Schlag (1966); Gilbert and Smith (1990).
II.2 Radiative transition rates
Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule Fermi (1932); Sakurai and Napolitano (2017)) and written in terms of Einstein coefficients in free space Einstein (1917); Cohen-Tannoudji, Diu, and Laloë (2019) and their generalizations in a weakly coupled and polaritonic microcavity. Rotational dynamics is much faster than thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
In free space, the spontaneous emission rate is given by the Einstein $A$ coefficientEinstein (1917); Cohen-Tannoudji, Diu, and Laloë (2019). In the case where spontaneous emission induces the $i\rightarrow j$ transition with $i>j$ this coefficient is given by
@@ II.2 Radiative transition rates
where $n_{\text{BE}}(\omega,T)$ is the Bose-Einstein distribution thermal occupation number $n_{\text{BE}}(\omega,T)=1/[\text{exp}(\hbar\omega/k_{B}T)-1]$Bose (1924) and $k_{B}$ is the Boltzmann constant.
The final pieces necessary to compute the rate constants above are the transition dipole matrix $\mu_{ij}$ and the frequencies $\omega_{ij}$. We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic molecules is given by the Morse potential [Fig. 1(a)]Morse (1929). The bound wave functions associated with the Morse potential are also employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
$\mu_{ij}$ $=\int\psi_{i}(r)\mu(r)\psi_{j}(r)\mathrm{d}r,$ (6)
where $\psi_{i}(r)$ represents the $i$th vibrational level wave function, and the dipole function $\mu(r)$ models the nonlinear change in the (electronic ground-state) molecular electrical dipole with the bond length $r$ Bishop (1990).
We investigated the IR radiative dissociation of hydrogen fluoride, HF, and lithium hydride, LiH. The employed dipole function $\mu(r)$ for HF is $\mu(r)=\mu_{0}re^{-\zeta r^{4}}$, where $\mu_{0}=0.45414e$ and $\zeta=0.081616$-1 were obtained from previous work on laser-induced HF dissociationKaluža et al. (1994). The LiH dipole function was interpolated from ab initio quantum chemical computations (see Supplementary Information). Morse potential parameters include the equilibrium bond length ($r_{e}$), electronic ground-state dissociation energy ($D_{e}$), and harmonic frequency ($\omega_{e}$) of each molecule and are listed in Table 1.
Table 1: Parameters used for Morse potential of HF Stine and Noid (1979); Guldberg and Billing (1991); Gross, Neuhauser, and Rabitz (1993); Kaluža et al. (1994) and LiH Irikura (2007).
@@
Note that the examined diatomic species have severely reduced room temperature BIRD rates relative to typical experiments involving polyatomic ions Dunbar and McMahon (1998); Dunbar et al. (1995) due to the relatively high stability and restricted number of state-space pathways available for diatomic dissociation. This feature motivated our choice of examining diatomic BIRD at temperatures $T=\mathcal{O}(\hbar\omega_{e}/k_{\text{B}})$, an order of magnitude greater than room temperature. Such temperatures lead to numerically stable BIRD rates in the various regimes considered in this work. Despite their largely different energetics, microcavity effects on HF and LiH BIRD yield the same qualitative behavior. This independence to microscopic details demonstrates the robustness of our analysis.
@@
$\omega_{\text{C}}(m,\textbf{q})=c\sqrt{q^{2}+\frac{m^{2}\pi^{2}}{L_{C}^{2}}}\;,$ (7)
where $\mathbf{q}=(q_{x},q_{y})\in\mathbb{R}^{2}$ is the in-plane projection of the mode wave vector, $q^{2}=q_{x}^{2}+q_{y}^{2}$, and $m$ is an integer quantum number that identifies the longitudinal wave vector $k_{z}=m\pi/L_{C}$ and thus the photonic energy bands in Fig. 1(c). For transverse-magnetic (TM) modes $m=0,1,2,...$ whereas $m=1,2,3...$ for transverse-electric (TE) modes Zoubi and La Rocca (2005); Jackson (2021). Including the TM0 mode here is crucial, as otherwise, transitions with energy below the cavity cutoff would be impossible.
Using Eq. 7, it can be shown (see Supporting Information) that the spatially-averaged electromagnetic DOS inside a microcavity in the weak coupling regime can be expressed by Barnes, Horsley, and Vos (2020)
@@
Lastly, we considered the BIRD scenario in which the diatomic interacts weakly with the electromagnetic component of polariton modes emergent from vibrational strong coupling between the microcavity modes and a material system (solid-state or molecular ensemble) with sufficiently large IR oscillator strength. We emphasize this material system does not undergo BIRD; its role is to generate polariton normal-modes with a tunable density of states and thermal energy density, as shown in Figs. 1(d) and (e). As we show in the SI, the formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
The Hamiltonian of the strongly coupled molecular subsystem and the microcavity is given in the Power-Zienau-Woolley (dipole) gauge Power and Zienau (1959); Woolley (1971)
@@
$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-m\pi c/L_{C}],$ (11)
where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the frequency-dependent photon content of polaritons with frequency $\omega$ (see SI)
@@
Our model shows singular behavior for $D_{P}(\omega)$ as $\omega\rightarrow\omega_{M}$. In this limit, both the photon content $P_{C}(\omega)$ and the group velocity $v_{g}(\omega,q)$ approach zero, but the ratio $P_{C}(\omega)/v_{g}(\omega,q)\rightarrow\infty$ resulting in $D_{P}(\omega)$ diverging to positive infinity. This singularity is due to the formation of the well-known stopgap regionLitinskaya and Agranovich (2009); Huang et al. (2003); Gómez Rivas et al. (2004); Todorov and Sirtori (2012) where $D_{P}(\omega)=0$. As previous work shows Satanin et al. (2005), introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$. Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on the type of disorder and coherence loss mechanism. This motivates our choice to work with the formally exact expression given in Eq. 11 and interpret our results as providing upper bounds for the effect of polaritons on BIRD.
@@ III Results and Discussion
III.1 Microcavity-assisted blackbody infrared radiative dissociation
Our analysis of radiative dissociation in weakly coupled microcavities assumes a single diatomic molecule (or a dilute sample at low enough pressure that collisions can be neglected) in a planar microcavity. The tunable electromagnetic environment afforded by the microcavity allows suggests BIRD rates can be either enhanced or suppressed relative to free space, depending on which transitions are significantly affected by the change induced by the microcavity in the photon DOS.
In Fig. 2, BIRD rates in the weak coupling regime are shown relative to the free space rate ($k_{0}=2.70\times 10^{-6}$ s-1 for HF at $T/\omega_{e}=0.67$ corresponding to $T=4000~{}\text{K}$ and $k_{0}=1.64\times 10^{-5}$ s-1 for LiH at $T/\omega_{e}=0.99$ corresponding to $T=2000~{}\text{K}$). For small microcavity lengths ($L_{C}<2$ $\mu$m for HF and $L_{C}<5$ $\mu$m for LiH), a consistent mild enhancement of the BIRD rate is associated with an overall increase in the microcavity electromagnetic DOS relative to free space over a wide frequency range covering most fundamental transitions and relevant overtones of HF and LiH. The ratios of microcavity BIRD rates to free space BIRD rates quickly approach 1.0 as $L_{C}$ increases beyond a few microns. This is expected since the microcavity DOS converges to the free space limit when $L_{C}\rightarrow\infty$.
The complex oscillatory pattern in the BIRD rates relative to free space shown in Fig. 2 arises from the interplay of transitions enhanced or suppressed depending on $L_{C}$. As we demonstrate below, this complexity is caused by overtones, which play an essential role in our analysis.
@@ III.1 Microcavity-assisted blackbody infrared radiative dissociation
To interpret the origin of the oscillatory behavior in Fig. 2 and identify the most consequential radiative transitions involved in BIRD of HF and LiH, we conducted a sensitivity analysis of the dissociation rate. This measures how impactful an enhancement in the photon DOS at a specific transition frequency is to the dissociation process (see Supporting Information for more details). For HF dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 23$, with $i\in\{16,17,19\}$ (in the relevant 1500 -- 5000 cm-1 frequency range) are the most consequential for the BIRD rate. Similarly, for LiH, the relevant transitions are $i\rightarrow 29$, with $i\in\{21,22,23,25\}$. In both cases, the identified overtones take the vibrational mode to the highest energy bound state (23 for HF and 29 for LiH) from which dissociation rapidly occurs. We emphasize that overtones are crucial because, at high excitation levels, fundamental transitions ($i\rightarrow i+1$) involve very small energies and, therefore, have associated small thermal photon populations and reduced oscillator strengths.
@@
Figure 3: Microcavity photon density of states normalized to the free space DOS for selected lengths ($L_{C}$) where BIRD suppression and enhancement were observed for HF (left panels) and LiH (right panels) in Fig. 2. The dotted vertical lines indicate the location of the transition frequencies corresponding to the most relevant overtone transitions, with color-coding to distinguish between suppression (red) and enhancement (green). Panel (a) shows the small HF BIRD suppression observed at $L_{C}=3.03~{}\mu\text{m}$ is at least in part caused by the reduced microcavity photon DOS relative to free space at the highlighted overtones. Panel (b) shows the mild enhancement of HF BIRD in a microcavity with $L_{C}=3.21~{}\mu\text{m}$ can be ascribed to the greater microcavity photon DOS at the frequencies of the highlighted overtones. Panels (c) and (d) show the analogous scenario for LiH, where BIRD is observed to be suppressed at $L_{C}=9.4~{}\mu\text{m}$ but enhanced at $L_{C}=9.6~{}\mu\text{m}$, primarily due to the variation of the microcavity photon density of states at the selected $i\rightarrow 29$ overtone frequencies.

### v2

@@ Abstract
Vibrational strong light-matter coupling offers a promising approach for controlling chemical reactivity with infrared microcavities. While recent research has examined potential mechanisms for this phenomenon, many important questions remain, including what type of reactions can be modified and to what extent this modification can be achieved. In this study, we explore the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes. Using a Master equation approach, we simulate the effects of infrared field confinement and vibrational strong coupling on BIRD rates for diatomic molecules. We present a framework explaining how infrared microcavities influence BIRD kinetics, highlighting the importance of overtone transitions in the process. Our findings outline the conditions under which significant enhancement or mild suppression of BIRD rates can be achieved, offering insights into the practical limitations and new strategies for controlling chemistry within infrared resonators.
@@ I   Introduction
Chemical reaction control is critical in applications ranging from industrial catalysis to medicine. Conventional methods for manipulating reaction rates involve varying temperature, pressure, or catalyst composition. An unconventional approach for controlling thermal reaction kinetics has been recently reported [1, 2, 3, 4] employing infrared (IR) microcavities [5] consisting of two moderate-quality mirrors separated by a distance $L_{C}$ of the order of IR wavelengths resonant with typical molecular vibrations [6]. When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device [7, 8, 9]. Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction rates[10, 11, 12, 13, 14, 15, 16, 4, 17, 18, 19]. Still, several fundamental questions remain [20, 21], including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted. In this study, we investigate gas-phase unimolecular dissociation activated by the absorption of microcavity blackbody infrared radiation [22] under weak and strong coupling with a suitable material.
@@
Blackbody infrared radiative dissociation (BIRD) was proposed by Perrin in 1913 as a mechanism for gas phase unimolecular dissociation [23]. However, experimental observations in the 1920s ruled BIRD out in favor of the collisional activation mechanism [24, 25, 26]. Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressures[27, 28, 29]. Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several simple and complex systems, including ion clusters, transition-metal complexes, and biopolymers [22].
@@
The thermal radiation spectrum heavily influences BIRD rates. Given that free space and microcavity electromagnetic density of states and thermal radiation density can be significantly different [5], microcavity BIRD is expected to proceed with different rates compared to free space BIRD. However, the extent of this change, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime remain open questions. Our study provides answers to these questions. Our quantitative analysis focuses on the dissociation of diatomic molecules induced by multiphoton or multipolariton absorption in microcavities, but several qualitative conclusions are generalizable to polyatomic systems.
@@
The primary aim of this study is to characterize thermal radiative dissociation in microcavities. We employ a Master equation treatment to investigate this process in (i) free space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material. Our results give upper bounds for BIRD enhancement and suppression in weakly and strongly coupled microcavities and identify overtone transitions as crucial ingredients for the observed modulation of BIRD. In Sec. II, we describe our methodology, Sec. III provides the main results and related discussion, and Sec. IV summarizes our main results and conclusions.
@@ II   Methods
Master Equation for diatomic BIRD
Our analysis of BIRD in microcavities employs similar tools to those successfully applied to investigate this process in free space [28, 30, 31, 22]. These include a Markovian Master equation describing the time-evolution of vibrational quantum state populations with energy below the dissociation threshold defined below. Rotational degrees of freedom are ignored since they are weakly coupled to vibrational transitions, and the rotational dynamics are essentially classical at the temperatures of interest to BIRD. We also neglect Doppler and lifetime broadening throughout. The impact of nonzero vibrational linewidths on the presented results is described in Results and Discussion.
For the sake of simplicity, we assumed the thermalization of the background radiation is faster than all considered radiative transitions. The rate of change of the population in the $i$th vibrational state ($N_{i}$) is given by
@@
where $k_{ij}$ denotes rate of absorption when $i>j$ and emission when $i<j$ and $i_{\text{max}}$ is the highest energy bound state (Fig. 1a). This state is only loosely bound and is assumed to undergo irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$. Equation 1 can be written in matrix notation as
@@
where N is a column vector containing the vibrational populations and $\mathbf{J}$ is the transition matrix. The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction [32, 33] of the described mechanism. Note that this mechanism does not consider dissociation from intermediate vibrational levels ($\nu<i_{\text{max}}$). Thus, our results are expected to be valid for BIRD promoted by multiphoton dissociation through a particular doorway state, which is the highest-energy bound state in our case.
Radiative transition rates
Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule [34, 35]) and written in terms of Einstein coefficients in free space [36, 37] and their generalizations in a weakly coupled and polaritonic microcavity. Rotational dynamics is much faster than thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
In free space, the spontaneous emission rate is given by the Einstein $A$ coefficient[36, 37]. In the case where spontaneous emission induces the $i\rightarrow j$ transition with $i>j$ this coefficient is given by
@@
and $n_{\text{BE}}(\omega,T)$ is the Bose-Einstein distribution thermal occupation number $n_{\text{BE}}(\omega,T)=1/[\text{exp}(\hbar\omega/k_{B}T)-1]$[38] and $k_{B}$ is the Boltzmann constant.
The final pieces necessary to compute the rate constants above are the transition dipole matrix $\mu_{ij}$ and the frequencies $\omega_{ij}$. We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic molecules is given by the Morse potential [Fig. 1(a)][39]. The bound wave functions associated with the Morse potential are also employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
$\mu_{ij}$ $=\int\psi^{*}_{i}(r)\mu(r)\psi_{j}(r)\mathrm{d}r,$ (6)
where $\psi_{i}(r)$ represents the $i$th vibrational level wave function, and the dipole function $\mu(r)$ models the nonlinear change in the (electronic ground-state) molecular electrical dipole with the bond length $r$ [40].
We investigated the IR radiative dissociation of hydrogen fluoride, HF, and lithium hydride, LiH. The employed dipole function $\mu(r)$ for HF is $\mu(r)=\mu_{0}re^{-\zeta r^{4}}$, where $\mu_{0}=0.45414e$ and $\zeta=0.081616$-4 were obtained from previous work on laser-induced HF dissociation[41]. The LiH dipole function was interpolated from ab initio quantum chemical computations (CCSD(T)/aug-cc-pVTZ, see Supplementary Information for further details). Morse potential parameters include the equilibrium bond length ($r_{e}$), electronic ground-state dissociation energy ($D_{e}$), and harmonic frequency ($\omega_{e}$) of each molecule and are listed in Table 1.
Table 1: Parameters used for Morse potential of HF [42, 43, 44, 41] and LiH [45].
@@
The examined dissociation processes (via the most energetic bound state) have reduced room temperature rates relative to typical BIRD experiments involving polyatomic ions [27, 46, 29, 22] due to the relatively high stability and restricted number of state-space pathways available for diatomic dissociation proceeding via the highest energy level. This feature motivated our choice of examining diatomic BIRD at temperatures $T=\mathcal{O}(\hbar\omega_{e}/k_{\text{B}})$, an order of magnitude greater than room temperature. The SI shows that the trends discussed below hold for a wide interval of temperatures. Despite their different energetics, microcavity effects on HF and LiH BIRD yield the same qualitative behavior.
@@
$\omega_{\text{C}}(m,q)=c\sqrt{q^{2}+\frac{m^{2}\pi^{2}}{L_{C}^{2}}}\;,$ (7)
where $\mathbf{q}=(q_{x},q_{y})\in\mathbb{R}^{2}$ is the in-plane projection of the mode wave vector, $q^{2}=q_{x}^{2}+q_{y}^{2}$, and $m$ is an integer quantum number that identifies the longitudinal wave vector $k_{z}=m\pi/L_{C}$ and thus the photonic energy bands in Fig. 1(c). For transverse-magnetic (TM) modes $m=0,1,2,...$ whereas $m=1,2,3...$ for transverse-electric (TE) modes [47, 48]. Including the TM0 mode here is crucial, as otherwise, transitions with energy below the cavity cutoff would be impossible.
Using Eq. 7, it can be shown (see Supporting Information) that the spatially-averaged electromagnetic DOS inside a microcavity in the weak coupling regime can be expressed by [49]
@@
Lastly, we considered the BIRD scenario in which the diatomic interacts weakly with the electromagnetic component of polariton modes emergent from vibrational strong coupling between the microcavity modes and a material system (solid-state or molecular ensemble) with sufficiently large IR oscillator strength. We emphasize this material system does not undergo BIRD; its role is to generate polariton normal-modes with a tunable density of states and thermal energy density, as shown in Figs. 1(d) and (e). The formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD (see SI) by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
The Hamiltonian of the strongly coupled molecular subsystem and the microcavity is given in the Power-Zienau-Woolley (dipole) gauge [50, 51]
@@
$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-\frac{m\pi c}{L_{C}}],$ (11)
where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the frequency-dependent photon content of polaritons with frequency $\omega$
@@
Our model shows singular behavior for $D_{P}(\omega)$ as $\omega\rightarrow\omega_{M}$. In this limit, both the photon content $P_{C}(\omega)$ and the group velocity $v_{g}(\omega,q)$ approach zero, but the ratio $P_{C}(\omega)/v_{g}(\omega,q)\rightarrow\infty$ resulting in $D_{P}(\omega)$ diverging to positive infinity. This singularity is due to the formation of the well-known stopgap region[59, 60, 61, 62] where $D_{P}(\omega)=0$. As previous work shows [63], introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$. Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on the type of disorder and coherence loss mechanism. This motivates our decision to use the formally exact expression presented in Eq. 11 and interpret our results as upper bounds on the effect of polaritons on BIRD.
@@ III   Results and Discussion
Microcavity-assisted blackbody infrared radiative dissociation
Our analysis of radiative dissociation in weakly coupled microcavities assumes a single diatomic molecule (or a dilute sample at low enough pressure that collisions can be neglected) in a planar microcavity. The tunable electromagnetic environment afforded by the microcavity allows BIRD rates to be either enhanced or suppressed relative to free space, depending on which transitions are significantly affected by the modified photon DOS.
In Fig. 2, BIRD rates in the weak coupling regime are shown relative to the free space rate ($k_{0}=2.70\times 10^{-6}$ s-1 for HF at $T/\omega_{e}=0.67$ corresponding to $T=4000~{}\text{K}$ and $k_{0}=1.64\times 10^{-5}$ s-1 for LiH at $T/\omega_{e}=0.99$ corresponding to $T=2000~{}\text{K}$). For small microcavity lengths, i.e., $L_{C}<2$ $\mu$m for HF and $L_{C}<5$ $\mu$m for LiH, a mild but consistent enhancement of BIRD rates is observed. At these lengths, an overall increase in the microcavity electromagnetic DOS over a wide frequency range covering several important transitions is associated with the observed enhancement. The ratios of microcavity BIRD rates to free space BIRD rates quickly approach 1.0 as $L_{C}$ increases beyond a few microns. This is expected since the microcavity DOS converges to the free space limit when $L_{C}\rightarrow\infty$.
The complex oscillatory pattern in the BIRD rates relative to free space shown in Fig. 2 arises from the interplay of discrete vibrational transitions enhanced or suppressed depending on $L_{C}$. As we demonstrate below, this complexity is caused by overtones, which play an essential role in our analysis.
@@
We conducted a sensitivity analysis of the dissociation rate to interpret the origin of the oscillatory behavior in Fig. 2 and identify the most consequential radiative transitions involved in the BIRD of HF and LiH. This analysis measures how impactful an enhancement in the photon DOS at a specific transition frequency is to the dissociation process (see Supporting Information for more details). For HF dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 23$, with $i\in\{16,17,19\}$ (in the relevant 1500 -- 5000 cm-1 frequency range) are the most consequential for the BIRD rate. Similarly, for LiH, the relevant transitions are $i\rightarrow 29$, with $i\in\{21,22,23,25\}$. In both cases, the identified overtones take the vibrational mode to the highest energy-bound state (23 for HF and 29 for LiH), from which dissociation rapidly occurs. We emphasize that overtones are crucial because, at high excitation levels, fundamental transitions ($i\rightarrow i+1$) involve very small energies and, therefore, have associated small thermal photon populations and reduced oscillator strengths. See Supporting Information for a brief discussion of results obtained without overtone transitions.
@@
Figure 3: Microcavity photon DOS normalized to the free space DOS for selected lengths ($L_{C}$) where BIRD suppression and enhancement were observed for HF (left panels) and LiH (right panels) in Fig. 2. The dotted vertical lines indicate the location of the transition frequencies corresponding to the most relevant overtone transitions, with color coding to distinguish between suppression (red) and enhancement (green). Panel (a) shows that the small HF BIRD suppression observed at $L_{C}=3.03~{}\mu\text{m}$ is at least in part caused by the reduced relative DOS at the highlighted overtones. Panel (b) shows the mild enhancement of HF BIRD in a microcavity with $L_{C}=3.21~{}\mu\text{m}$ can be ascribed to the greater microcavity photon DOS at the frequencies of the highlighted overtones. Panels (c) and (d) show the analogous scenario for LiH, where BIRD is observed to be suppressed at $L_{C}=9.4~{}\mu\text{m}$ but enhanced at $L_{C}=9.6~{}\mu\text{m}$, primarily due to the variation of the microcavity photon density of states at the selected $i\rightarrow 29$ overtone frequencies.